# Multipole moments do not uniquely characterise spacetimes beyond general relativity

### v1

@@ Abstract
Spacetimes in general relativity can be uniquely decomposed into a set of multipole moments. Given the usefulness of moments in the categorisation of radiation patterns, tidal deformations, and other phenomena associated with compact objects, a number of studies have explored their construction in beyond-Einstein theories of gravity. It is shown here that uniqueness does not necessarily extend across theories: by comparing a few static and spherically-symmetric solutions in different theories, we find that two distinct objects can possess the same Geroch-Hansen moments. Moreover, two metrics can match and yet take different moments. Implications of this result are explored in the context of black-hole shadows and “universal” relations hinging on moment computations.
@@ I Introduction
Although the procedure for computing moments is not unique in GR, it has been established that a given set of moments uniquely corresponds to a spacetime and the prescriptions considered in the literature are equivalent within the appropriate realms of validity Xanthopoulos (1979); Beig and Simon (1981); Simon and Beig (1983). Notably, Gürsel (1983) showed that the Geroch-Hansen Geroch (1970a, b); Hansen (1974) moments are equivalent to the Thorne (1980) moments for stationary systems (see also Ref. van Voorthuizen (2025) for some refinements with respect to proof gaps). However, the goal of this paper is to demonstrate that dropping the qualifier “in GR” renders the more-general statement untrue.
We consider a few example, asymptotically-flat solutions in some scalar-tensor theories of gravity and compute their moments using a Geroch-Hansen procedure. While the line elements and scalar profiles differ, it is shown that their moments can match; this implies that knowledge of the latter is insufficient to reconstruct a spacetime in the absence of a fixed theory. This result appears to have gone unpenned in the literature, though could have important consequences. For example, “universal relations” involving quadrupole moments and other intrinsic parameters of compact objects have been extensively explored in a variety of contexts Yagi and Yunes (2013); Pappas and Apostolatos (2014); Yagi et al. (2014, 2014); Martinon et al. (2014); Pappas et al. (2019). If, however, the moments themselves vary across theories, it becomes less clear how one may utilise these relations unless one a priori assumes a fixed theory of gravity when comparing predictions to observational data.

### v2

# Multipole moments do not uniquely characterize spacetimes beyond general relativity

@@ Abstract
Spacetimes in general relativity can be uniquely decomposed into a set of multipole moments. Given the usefulness of moments in the categorization of radiation patterns, tidal deformations, and other phenomena associated with compact objects, a number of studies have explored their construction in beyond-Einstein theories of gravity. It is shown here that uniqueness does not necessarily extend across theories: by comparing a few static and spherically-symmetric solutions in different theories, we find that two distinct objects can possess the same Geroch-Hansen moments. Moreover, two metrics can match and yet take different moments. Implications of this result are explored in the context of black-hole shadows and “universal” relations hinging on moment computations.
@@ I Introduction
Although the procedure for computing moments is not unique in GR, it has been established that a given set of moments determines a vacuum metric up to isometries and the prescriptions considered in the literature are equivalent within the appropriate realms of validity Xanthopoulos (1979); Beig and Simon (1981); Simon and Beig (1983); Beig and Schmidt (2000). Notably, Gürsel (1983) showed that the Geroch-Hansen moments are equivalent to the Thorne moments for stationary systems (see also Ref. van Voorthuizen (2025) for some refinements with respect to proof gaps). However, the goal of this paper is to demonstrate that dropping the qualifier “in GR” renders the more-general statement untrue.
We consider a few example, asymptotically-flat solutions in some scalar-tensor theories of gravity and compute their moments using a Geroch-Hansen procedure. While the line elements and scalar profiles differ, it is shown that their moments can match; this implies that knowledge of the latter is insufficient to reconstruct a spacetime in the absence of a fixed theory. This result appears to have gone unpenned in the literature, though could have important consequences. For example, “universal relations” involving quadrupole moments and other intrinsic parameters of compact objects have been extensively explored in a variety of contexts Yagi and Yunes (2013); Pappas and Apostolatos (2014); Yagi et al. (2014, 2014); Martinon et al. (2014); Pappas et al. (2019); Adam et al. (2024). If, however, the moments themselves vary across theories, it becomes less clear how one may utilise these relations unless one a priori assumes a fixed theory of gravity when comparing predictions to observational data. We emphasise at this point that we do not address the harder problem of when a given set of moments uniquely corresponds to solutions within a fixed theory. Such a problem boils down to assessing whether Birkhoff-like theorems hold, which generally requires case-by-case handling Ghosh et al. (2025); some first steps towards a result are presented in Appendix A.
@@ Acknowledgements.
Appendix A Non-uniqueness within a fixed theory
In addition to the results discussed in the main text, it is worth pointing out that even within a given theory a set of moments may fail to uniquely capture vacuum structure. This can be demonstrated – again by counterexample – by simply noting that any free parameters appearing within the action of a theory will typically manifest within solutions. As such, since only monopole moments can be non-zero in static and spherically-symmetric cases, if the (uncharged) theory hosts at least two coupling parameters then it is clear that the two available moments cannot contain enough information to catalogue solutions (i.e., the mass and two innate hairs).
Such a restriction is not necessary however, as even some theories with only a single, short-range parameter may be degenerate in the sense of moments. In fact, even some theories with zero free action parameters may be degenerate. One obvious example is the pure $R^{2}$ theory. Any metric with vanishing Ricci scalar is a solution in these theories Suvorov (2021) and thus, since there are (at least) two free metric functions $A$ and $B$ in expression (1), an arbitrary number of free parameters can appear in a static solution. A small set of monopole moments clearly cannot encode such near-field details in general.
As a less contrived example, consider the hybrid Gauss-Bonnet-Horndeski theory introduced by Lü and Pang (2020), whose action reads
$S=$ $\int d^{4}x\sqrt{-g}\Big\{R+\alpha\big[\phi\mathcal{G}+4G^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi$ (29)
$-2\chi Re^{-2\phi}-4\nabla_{\mu}\phi\nabla^{\mu}\phi(\nabla_{\nu}\nabla^{\nu}\phi)$
[…]
$-6\chi^{2}e^{-4\phi}\big]\Big\}.$
Here, $G^{\mu\nu}$ is the Einstein tensor, $\mathcal{G}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet invariant, and GR is recovered by setting the coupling parameter, $\alpha$, to zero. The other parameter, $\chi$, is inessential and can effectively be absorbed by field redefinitions Babichev et al. (2023). There exists an exact solution in this theory with metric taking the form of expression (1) with $B=1/A$ and $D=r^{2}$, such that
$A(r)$ $=1+\frac{r^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r^{3}}}\right)$ (30)
$=1-\frac{2M}{r}+\frac{4\alpha M^{2}}{r^{4}}+\mathcal{O}\left(\frac{1}{r^{7}}\right).$
The associated scalar field can be written in full in terms of integrals of $A$, though has the convenient expansion
$\phi(r)=\frac{M}{r}+\frac{3M^{2}}{4r^{2}}+\frac{5M^{3}}{6r^{3}}+\mathcal{O}\left(\frac{1}{r^{4}}\right).$ (31)
What is interesting is that the parameter $\alpha$ appears only at order $1/r^{4}$ in the potential (30) and not at all within the scalar field (31). Though not shown rigorously here, it is clear that the mass- and scalar-monopole moments for this solution thus both take value $M$ (up to sign),
$M_{0}=M,\qquad\Phi_{0}=M.$ (32)
Since the parameter $\alpha$ does not appear in expressions (32), this implies that there are infinitely many solutions within the theory described by (29) corresponding to a given set of moments. Care must be exercised therefore in describing “universal relations” involving multipole moments within this or similar theories.